# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 6.1. Admissible space of mixed discrete conformal structure $\mathrm{II}$
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,-1)$. Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}>-1$ and $(u_{i}+u_{j})\in\mathbb{R}$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$. Since $u_{i}>0,u_{j}<0$, then $\mathrm{arcsinh}\ (-\xi_{jk})<u_{j}+u_{k}<u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$ and hence $\xi_{ik}+\xi_{jk}>0$. Then
$\mathcal{U}^{(14)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>-1,\ \xi_{ik}\in\mathbb{R}$ and $\xi_{ik}+\xi_{jk}>0$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,-1)$. Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$. Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$. Then $\mathrm{arcsinh}\ (-\xi_{jk})+\mathrm{arcsinh}(-\xi_{ij})<u_{i}+2u_{j}+u_{k}<2u_{j}+\mathrm{arccosh}\xi_{ik}<\mathrm{arccosh}\xi_{ik}$. For simplicity, we require $\xi_{ij}>0$. Then
$\mathcal{U}^{(15)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij}),$
$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>1$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,-1,-1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$ in Subsection 3.1. Moreover, $\xi_{ij}>0,\ u_{i}+u_{j}<\log\xi_{ij}$ and $\xi_{ik}>0,\ u_{i}+u_{k}<\log\xi_{ik}$. Then
$\mathcal{U}^{(16)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{j}<\log\xi_{ij},$
$u_{i}+u_{k}<\log\xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}>0,\ \xi_{ik}>0$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,-1,-1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$. Since $u_{i}>0$ and $u_{k}<0$, then $-\mathrm{arccosh}\ \xi_{jk}<u_{j}+u_{k}<u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$. For simplicity, we assume $\xi_{ij}>0$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$, then we assume $\xi_{ik}>0$ for simplicity. Then
@@
$u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij},\ u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}>0,\ \xi_{ik}>0$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,-1,-1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}>1$ and $-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij}$. Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$. Then
@@
$-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij},$
$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}>1,\ \xi_{ik}>1$.
In the above different admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$, the weight $\xi$ has different range of values. We are unable to use a uniform formula to express them. For simplicity, we use a notation $\Xi$ to denote the set of the range of weight $\xi$ in the whole admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$.
@@
Suppose $(\Sigma,\mathcal{T},\alpha,\xi)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\xi\in\Xi$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (46). Then the admissible space $\mathcal{U}^{(n)}_{ijk}(\xi)$ of $u$ has 18 types, each of which is convex polytope. As a result, the admissible space
$\mathcal{U}(\xi)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\xi),\ \mathcal{U}_{ijk}(\xi)\in\{\mathcal{U}^{(n)}_{ijk}(\xi),\ n=1,2,...,18\}$
is also a convex polytope.
Remark 6.2.
The admissible space $\mathcal{U}(\xi)$ is a finite intersection of 18 types of admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$ for $n=1,2,...,18$. But this finite intersection is not arbitrary because of the connectedness of the admissible space $\mathcal{U}(\xi)$. For example, if there exist the admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$ and $\mathcal{U}^{(18)}_{ijk}(\xi)$ for some right-angled hyperbolic hexagons $\{ijk\}\in F$ with edge lengths given by (46), then there must be the third type of the admissible space $\mathcal{U}^{(n)}_{ijk}(\xi)$ for $n\neq 1,18$.
6.1.1. Negative definiteness of Jacobian
@@
Under the same assumptions as in Theorem 6.1. The Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite. As a consequence, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on the admissible space $\mathcal{U}(\xi)$.
@@
For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. By (47), we have
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{6},$
where $C_{r}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$. Then
where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Note that $d_{ij}<0,d_{ik}<0$ and the others are positive for any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46). The rest of the proof is similar to that of Theorem 5.2, so we omit it. Q.E.D.
6.2. Rigidity of the mixed discrete conformal structure $\mathrm{II}$
The following theorem gives the rigidity of mixed discrete conformal structure $\mathrm{II}$, which generalizes the case $\mathrm{(iv)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it.
@@ 6.2. Rigidity of the mixed discrete conformal structure $\mathrm{II}$
Under the same assumptions as in Theorem 6.1. The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{E}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding.
6.3. Existence of the mixed discrete conformal structure $\mathrm{II}$
We consider the image of $K$ for four types of the admissible spaces $(0,0,0),(1,0,0),(0,0,1),(1,0,1)$. In other words, we do not consider the case that $\alpha_{j}=\alpha_{k}=1$ for $\alpha\in\{0,1\}$. For simplicity, we use a new notation $\widetilde{\Xi}$ to denote the set of the range of weights $\xi$ in the admissible spaces $\mathcal{U}^{(n)}_{ijk}$, i.e., $n=1,2,4,5.$
@@ 6.3. Existence of the mixed discrete conformal structure $\mathrm{II}$
Suppose $(\Sigma,\mathcal{T},\alpha,\xi)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{0,1\}$ and $\xi\in\widetilde{\Xi}$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (46). If $\alpha_{j}$ and $\alpha_{k}$ can not take 1 simultaneously, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
@@
By (46), we have
(56) $\cosh l_{ij}=\bigg{[}\xi_{ij}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
$\cosh l_{ik}=\bigg{[}\xi_{ik}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{i}+f_{k}},$
$\cosh l_{jk}=\bigg{[}\xi_{jk}-\sqrt{(e^{-2f_{j}}+\alpha_{j})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{j}+f_{k}}.$
For the case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. Then $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{i}+f_{j}})$$\lim\cosh l_{ik}:=\lim(c_{j}e^{f_{i}+f_{k}})$ and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants. By the cosine law,
@@
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs of the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$ are similar, so we omit them here.
Q.E.D.
@@
Under the same assumptions as in Theorem 6.5. If $\lim f_{k}=-\infty$, then $\lim f_{j}=+\infty$. Moreover, if one of the following conditions is satisfied
@@
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$;
@@
Combining (47) and Remark 2.23, if $\lim f_{k}=-\infty$, then $\lim u_{k}=-\infty$. Note that $l_{jk}>0$ implies that $u_{j}+u_{k}>C(\xi_{jk})$, where $C(\xi_{jk})$ is a constant that depends on $\xi_{jk}$. Then $\lim u_{j}=+\infty$. This implies $\alpha_{j}=0$ and $\lim f_{j}=+\infty$ by Remark 2.23.
Set $\widetilde{L}_{1}=\frac{\cosh l_{ij}}{\sinh l_{ik}\sinh l_{jk}}$ and $\widetilde{L}_{2}=\frac{\cosh l_{ik}\cosh l_{jk}}{\sinh l_{ik}\sinh l_{jk}}$, then $\cosh\theta_{k}^{ij}=\widetilde{L}_{1}+\widetilde{L}_{2}$. Moreover, (46) is reduced to
$\cosh l_{ij}$ $=\sqrt{1+\alpha_{i}e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{i}e^{2f_{i}}}+\xi_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $\rightarrow-1+\xi_{jk}e^{f_{j}+f_{k}}.$
For the case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}=\lim(-1+\xi_{jk}e^{f_{j}+f_{k}})$ and
$\lim\cosh l_{ik}=\begin{cases}\lim(1+\xi_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{i}=0,\\ \lim(c_{j}e^{f_{i}}),\ &\text{if}\ \alpha_{i}=1,\end{cases}$

### v3

@@ 8.1. Admissible space of the mixed discrete conformal structure I
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,-1)$. It is known that $u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk})$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}>-1$ and $(u_{i}+u_{j})\in\mathbb{R}$. Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$. Since $u_{i}>0$ and $u_{j}<0$, it follows that $\mathrm{arcsinh}\,(-\eta_{jk})<u_{j}+u_{k}<u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$, and hence $\eta_{ik}+\eta_{jk}>0$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(14)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}\},$
where $\eta_{jk}>0$, $\eta_{ij}>-1$, $\eta_{ik}\in\mathbb{R}$, and $\eta_{ik}+\eta_{jk}>0$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,-1)$. It is known that $u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk})$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij})$. Similarly, $\eta_{ik}>1$ and $-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}$. Combining these results, we derive the inequality chain: $\mathrm{arcsinh}\,(-\eta_{jk})+\mathrm{arcsinh}\,(-\eta_{ij})<u_{i}+2u_{j}+u_{k}<2u_{j}+\mathrm{arccosh}\,\eta_{ik}<\mathrm{arccosh}\,\eta_{ik}$. For simplicity, we assume $\eta_{ij}>0$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(15)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij}),$
$\ -\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>1$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,-1,-1)$. In Subsection 3.1, it follows that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$. Moreover, $\eta_{ij}>0$ with $u_{i}+u_{j}<\log\eta_{ij}$, and $\eta_{ik}>0$ with $u_{i}+u_{k}<\log\eta_{ik}$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(16)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{j}<\log\eta_{ij},$
$\ u_{i}+u_{k}<\log\eta_{ik}\},$
where $\eta_{jk}>1$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,-1,-1)$. It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$. Moreover, $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij}$. Since $u_{i}>0$ and $u_{k}<0$, it follows that $-\mathrm{arccosh}\,\eta_{jk}<u_{j}+u_{k}<u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij}$. For simplicity, we assume $\eta_{ij}>0$. Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$, thus we assume $\eta_{ik}>0$. Consequently, the admissible space of $u$ is defined as
@@
$\ u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij},\ u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}\},$
where $\eta_{jk}>1$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,-1,-1)$. It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$. Moreover, $\eta_{ij}>1$ and $-\mathrm{arccosh}\,\eta_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\,\eta_{ij}$. Similarly, $\eta_{ik}>1$ and $-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}$. Consequently, the admissible space of $u$ is defined as
@@
$-\mathrm{arccosh}\,\eta_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\,\eta_{ij},$
$-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}\},$
where $\eta_{jk}>1$, $\eta_{ij}>1$, and $\eta_{ik}>1$.
Across the various admissible spaces $\mathcal{U}^{(n)}_{ijk}(\eta)$ discussed above, the weight $\eta$ takes on different value ranges. These ranges cannot be expressed through a single consistent formula. For simplicity, we introduce the notation $\Xi$ to denote the set of all such weight ranges associated with the entire family of admissible spaces $\mathcal{U}^{(n)}_{ijk}(\eta)$, where $n=1,2,\dots,18$.
@@
Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (53). Then the admissible space $\mathcal{U}^{(n)}_{ijk}(\eta)$ of $u$ admits 18 distinct types, each of which is a convex polytope.
We rewrite the admissible space in (35) as $\mathcal{U}^{(0)}_{ijk}(\eta)$. Combining Theorem 3.1 and Theorem 8.1 gives the following result.
Corollary 8.2.
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where $\alpha:B\rightarrow\{-1,0,1\}$. If one of the following conditions holds:
(1):
All right-angled hyperbolic hexagons are Type-I (i.e., the edge lengths of any right-angled hyperbolic hexagons are given by (53)), and the weight $\eta$ belongs to $\Xi$;
(2):
Not all right-angled hyperbolic hexagons are Type-I (i.e., there exist some right-angled hyperbolic hexagons with edge lengths given by (5)). In this case, for those right-angled hyperbolic hexagons with edge lengths given by (53), the weight $\eta$ belongs to $\Xi$, while for those with edge lengths given by (5), $\eta$ satisfies the condition in Theorem 3.1.
Then the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is a convex polytope on $(\Sigma,\mathcal{T},\eta)$. Here $\mathcal{U}_{ijk}(\eta)\in\{\mathcal{U}^{(n)}_{ijk}(\eta),\ n=0,1,2,...,18\}$.
8.2. Negative definiteness of the Jacobian
@@
Under the same assumptions as those in Theorem 8.1, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite on $\mathcal{U}_{ijk}(\eta)$.
@@
For simplicity, we adopt the notations introduction in Subsection 3.2. Formula (54) implies
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{8},$
where $C_{r}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$. Then $\Lambda_{ijk}$ can be expressed as
$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{8},$
where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively. Furthermore, $\det Q_{1}>0$. For any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (53), since $d_{ij}<0$, $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$, it follows from (39) that $\det Q_{2}<0$. The remaining part of the proof is analogously to that of Theorem 6.3, and is thus omitted here.
Q.E.D.
Corollary 8.4.
Under the same assumptions as those in Corollary 8.2, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
8.3. Rigidity of the mixed discrete conformal structure I
The following theorem establishes the rigidity of the mixed discrete conformal structure I, which generalizes the rigidity part of Theorem 1.8 (iv). As its proof is similar to that of Theorem 3.4, it is omitted here.
@@ 8.3. Rigidity of the mixed discrete conformal structure I
Under the same assumptions as those in Corollary 8.2, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
8.4. Existence of the mixed discrete conformal structure I
For Type-I right-angled hyperbolic hexagons, we restrict our consideration to four types of admissible spaces, specifically corresponding to $(0,0,0)$, $(1,0,0)$, $(0,0,1)$, and $(1,0,1)$. These restrictions are motivated by two considerations: first, all cases where $(\alpha_{i},\alpha_{j},\alpha_{k})$ includes $-1$ are excluded, as the inclusion of $-1$ introduces significant technical complexities in the proof. Second, the case $\alpha_{j}=\alpha_{k}=1$ is also excluded, with a detailed explanation provided in Remark 8.10. For simplicity, we introduce a new notation $\widetilde{\Xi}$ to denote the set consisting of the ranges of weights $\eta$ across all admissible spaces $\mathcal{U}^{(n)}_{ijk}$ with $n=1,2,4,5$.
The following theorem characterizes the image of $K$ for the mixed discrete conformal structure I, which generalizes the existence part of Theorem 1.8 (iv).
@@ 8.4. Existence of the mixed discrete conformal structure I
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where $\alpha:B\rightarrow\{0,1\}$. If one of the following conditions holds:
(1):
All right-angled hyperbolic hexagons are Type-I (i.e., the edge lengths of any right-angled hyperbolic hexagon are given by (53)), $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously, and $\eta\in\widetilde{\Xi}$;
(2):
Not all right-angled hyperbolic hexagons are Type-I (i.e., there exist some right-angled hyperbolic hexagons with edge lengths given by (5)). In this case, for those right-angled hyperbolic hexagons with edge lengths given by (53), $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously, and $\eta\in\widetilde{\Xi}$, while for those with edge lengths given by (5), $\eta$ satisfies the condition in Theorem 3.1.
Then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weights are given by $\alpha:B\rightarrow\{0,1\}$ and $\eta\in\widetilde{\Xi}$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (53), and assume $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously. If one of the following conditions is satisfied
@@
Formula (53) implies
(63) $\cosh l_{ij}=\bigg{[}\eta_{ij}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
$\cosh l_{ik}=\bigg{[}\eta_{ik}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{i}+f_{k}},$
$\cosh l_{jk}=\bigg{[}\eta_{jk}-\sqrt{(e^{-2f_{j}}+\alpha_{j})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{j}+f_{k}}.$
For the case (i), it follows that $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{i}+f_{j}})$, $\lim\cosh l_{ik}:=\lim(c_{j}e^{f_{i}+f_{k}})$, and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law, we have
@@
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs for the cases (ii) and (iii) follow analogously and are omitted here. Q.E.D.
@@
Under the same assumptions as those in Lemma 8.7, if $\lim f_{k}=-\infty$, then $\lim f_{j}=+\infty$. Moreover, if one of the following conditions is satisfied
@@
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$,
@@
Combining (54) and Remark 2.22, if $\lim f_{k}=-\infty$, then $\lim u_{k}=-\infty$. As established in Subsection 8.1, the condition $l_{jk}>0$ implies $u_{j}+u_{k}>C(\eta_{jk})$, where $C(\eta_{jk})$ is a constant depending on $\eta_{jk}$. This, in turn, forces $\lim u_{j}=+\infty$. By combining (54) and Remark 2.22 again, we have $\alpha_{j}=0$ and $\lim f_{j}=+\infty$.
Set $\widetilde{L}_{1}=\frac{\cosh l_{ij}}{\sinh l_{ik}\sinh l_{jk}}$ and $\widetilde{L}_{2}=\frac{\cosh l_{ik}\cosh l_{jk}}{\sinh l_{ik}\sinh l_{jk}}$, then $\cosh\theta_{k}^{ij}=\widetilde{L}_{1}+\widetilde{L}_{2}$. Moreover, (53) reduces to
$\cosh l_{ij}$ $=\sqrt{1+\alpha_{i}e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{i}e^{2f_{i}}}+\eta_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $\rightarrow-1+\eta_{jk}e^{f_{j}+f_{k}}.$
For the case (i), it follows that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$, $\lim\cosh l_{jk}=\lim(-1+\eta_{jk}e^{f_{j}+f_{k}})$, and
$\lim\cosh l_{ik}=\begin{cases}\lim(1+\eta_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{i}=0,\\ \lim(c_{j}e^{f_{i}}),\ &\text{if}\ \alpha_{i}=1,\end{cases}$